# New phenomena arising from ${\mathcal{L}}$-invariants of modular forms

### v1

@@ 9.1. High multiplicity slopes
To answer the first question, we performed the following experiment. For each $v$, we gathered the total multiplicity of $v_{p}({\mathcal{L}})$ occurring in the closed intervals $[v-2,v+2]$. While $2$ is somewhat arbitrary, recording multiplicities over intervals of positive length captures clustering we see in our figures, for instance around around $v_{p}({\mathcal{L}})=-10,-11,-12$ in Figure 6. The results for levels $\Gamma_{0}(3)$, $\Gamma_{0}(5)$, and $\Gamma_{0}(7)$ are displayed, respectively, in Figures 21, 22, and 23.
@@ 9.3. On the integrality of slopes of ${\mathcal{L}}$-invariants
In fact, this global phenomenon is a consequence of a local conjecture. Namely, assume $v_{p}(a)>0$ and let $r_{k,a}$ denote the unique irreducible crystalline representation with Hodge-Tate weights 00 and $k-1$ on which Frobenius has characteristic polynomial $x^{2}-ax+p^{k-1}$. The local conjecture proposes that when $\overline{r}_{k,a}$ is regular and $k$ is even, then $v_{p}(a)$ is integral. See [21, Conjecture 4.1.1]. Under relatively mild hypotheses, this local conjecture (and thus the global one) has been proven in [52, Theorem 1.9] using the machinery of the ghost conjecture.

### v2

@@ 9.1. High multiplicity slopes
To answer the first question, we performed the following experiment. For each $v$, we gathered the total multiplicity of $v_{p}({\mathcal{L}})$ occurring in the closed intervals [v2,v+2][v-2,v+2]. While $2$ is somewhat arbitrary, recording multiplicities over intervals of positive length captures clustering we see in our figures, for instance around around $v_{p}({\mathcal{L}})=-10,-11,-12$ in Figure 6. The results for levels $\Gamma_{0}(3)$, $\Gamma_{0}(5)$, and $\Gamma_{0}(7)$ are displayed, respectively, in Figures 21, 22, and 23.
@@ 9.3. On the integrality of slopes of $\mathcal{L}$}-invariants
In fact, this global phenomenon is a consequence of a local conjecture. Namely, assume $v_{p}(a)>0$ and let $r_{k,a}$ denote the unique irreducible crystalline representation with Hodge-Tate weights 0 and k1k-1 on which Frobenius has characteristic polynomial $x^{2}-ax+p^{k-1}$. The local conjecture proposes that when $\overline{r}_{k,a}$ is regular and $k$ is even, then $v_{p}(a)$ is integral. See [21, Conjecture 4.1.1]. Under relatively mild hypotheses, this local conjecture (and thus the global one) has been proven in [51, Theorem 1.9] using the machinery of the ghost conjecture.